A balloon is being inflated by pumped air at the rate of 2 cubic inches per second. How fast is the diameter of the balloon increasing when the radius is ½ inch?


1
Expert's answer
2021-05-24T16:23:25-0400

Volume of spherical balloon is given by, V = "\dfrac{4{\pi}r^{3}}{3}" ,

Now differentiating with respect to r we get ,


"\dfrac{dV}{dr}=4{\pi}r^{2}"


"dV=4{\pi}r^{2}dr" .....1)


"\dfrac{dV}{dt}=2" cubic inches per second


"\dfrac{dr}{dt}="


Radius r = "\dfrac{1}{2}" inch



now differentiating equation 1 with respect to time , we get "\dfrac{dV}{dt}=4{\pi}r^{2}\dfrac{dr}{dt}"


now puting all the value in the given equation , we get -


= "2=" "4{\pi}\dfrac{1}{4}
\dfrac{dr}{dt}"


="\dfrac{dr}{dt}=\dfrac{2}{\pi}" inches ......2) , relation between diameter and radius is given by r = "\dfrac{D}{2}"


="\dfrac{d(D)}{dt}=\dfrac{4}{\pi}\ inches"